400 tickets sold, students $2.00, adults $3.00. Took in $1050.00 How many of each sold?
I think the equation is $2.00x+$3.00=$1050.00
x+y=400

Answer by Nate(3500)

Correct, except you are missing a variable: x = child y = adult
2.00x + 3.00y = 1050.00
2x + 3(400 - x) = 1050
2x + 1200 - 3x = 1050
-x = -150
x = 150
plug:
400 - x = y
400 - 150 = y = 250
250 adult's tickets and 150 children's tickets